OpenStudy (jaedyn):

Solve for x: one over four (4x + 15) = 24 49 two hundred seven over four eighty one over four 9

OpenStudy (jaedyn):

Help

OpenStudy (davejavous):

Re-writing your question: \[ \frac{1}{4(4x + 15)} = 24 \] (a) Firstly, expand the bracket on the bottom of the fraction. (b) Secondly, get rid of the fraction by multiplying both sides with what you got from part (a).

OpenStudy (will.h):

Is that what you mean? 1/4( 4x + 15) = 24 If yes Distribute 1/4 X + 3.75 = 24 Subtract 3.75 from both sides X = 24 - 3.75 X = 20.25

OpenStudy (will.h):

By the way x = 20.25 Is equivalent to eight one over four

OpenStudy (will.h):

Eighty one over four *